a pole support bby steel wire is in the way of construction site.in order to move heavy machines the steel wire must be moved 4m closer to the pole if the pole height is 30m and cable is currently fastened to the ground 12m from the pole how much steel wire need to cut off from the cable when they move it
1
Expert's answer
2019-10-11T10:47:37-0400

Initial length of wire:


"l_0=\sqrt{30^2+12^2}=32.311\ m"

Finial length of wire:


"l=\sqrt{30^2+(12-4)^2}=31.048\ m"

"\Delta l=32.311-31.048=1.263\ m"